# Addressing the Rank Degeneration in Sequential Recommendation via Singular Spectrum Smoothing

### v1

@@ 6. Conclusion
We investigate the rank degeneration issue in both sequence output embeddings and item embeddings, which are observed as the fast singular value decay phenomenon. We argue that the rank degeneration issues in both sequences and items are intertwined. With theoretical demonstrations, we show the possibility of sequence rank collapse in SR, which is more severe in short sequence learning. We propose a novel metric to measure the degree of rank degeneration, which is calculated as the area under the singular value curve (AUSC). We further propose a singular spectrum smoothing regularization as a surrogate of AUSC to constrain the Transformer and item embeddings learning, alleviating the rank collapse issue and preserving the ranks of embeddings. We also connect the ranks of sequence and item embeddings with the rank of the user-item preference prediction matrix, which correlates with the recommendation diversity. We empirically demonstrate the effectiveness of the proposed singular spectrum smoothing regularization in recommendation performances, especially for short sequences and popular items. We also observe a strong connection between the proposed singular spectrum smoothing regularization and the recommendation diversity.

### v2

@@ 7. Conclusion
We theoretically and empirically identify the relationship between the representation degeneration and recommendation diversity. Motivated by this connection, we propose a novel metric to measure the degree of degeneration as the area under the singular value curve (AUSC) to constrain the sequence and item embeddings learning. We further propose a singular spectrum smoothing regularization as a surrogate of AUSC to alleviate the degeneration and improve the diversity simultaneously. We empirically demonstrate the effectiveness of the proposed singular spectrum smoothing regularization in recommendation performances, especially for short sequences and popular items. We also empirically demonstrate a strong connection between the proposed singular spectrum smoothing regularization and the recommendation diversity.
Appendix A Baselines and Hyper-parameters Grid Search
We implement SPMRec with Pytorch based on the S3-Rec (Zhou et al., 2020) code base. We grid search all parameters and report the test performance based on the best validation results. We search the representation dimension in $\{64,128\}$, maximum sequence length from $\{50,100\}$, learning rate in $\{10^{-3},10^{-4}\}$, the $L2$ regularization control weight from $\{0.0,10^{-1},10^{-2},10^{-3}\}$, dropout rate from $\{0.3,0.5,0.7\}$ for all methods. For sequential methods, we search number of layers from $\{1,2,3\}$, and number of heads in $\{1,2,4\}$. We adopt the early stopping strategy that model optimization stops when the validation MRR does not increase for 50 epochs. The followings are the model specific hyper-parameters search ranges of baselines:
  • BPR [^2] https://github.com/xiangwang1223/neural_graph_collaborative_filtering: BPR is the most classical collaborative filtering method for personalized ranking with implicit feedbacks. We search the learning rate in $\{10^{-3},10^{-4}\}$ and $L2$ regularization weight from $\{10^{-1},10^{-2},10^{-3}\}$.
  • Caser [^3] https://github.com/graytowne/caser_pytorch: A CNN-based sequential recommendation method that views the sequence embedding matrix as an image and applies convolution operators to it. We search the length $L$ from $\{5,10\}$, and $T$ from $\{1,3,5\}$.
  • SASRec [^4] https://github.com/RUCAIBox/CIKM2020-S3Rec: The state-of-the-art sequential method that depends on the Transformer architecture. We search the dropout rate from $\{0.3,0.5,0.7\}$.
  • BERT4Rec [^5] https://github.com/FeiSun/BERT4Rec: This method extends SASRec to model bidirectional item transitions with standard Cloze objective. We search the mask probability from the range of $\{0.1,0.2,0.3,0.5,0.7\}$.
  • DT4SR [^6] https://github.com/DyGRec/DT4SR: A metric learning-base sequential method that models items as distributions and proposes mean and covariance Transformers. We search the dropout rate from $\{0.3,0.5,0.7\}$.
  • STOSA [^7] https://github.com/zfan20/STOSA: A metric learning-base sequential method that models items as distributions and proposes a Wasserstein self-attention module. We search the dropout rate from $\{0.3,0.5,0.7\}$.
  • DuoRec: [^8] https://github.com/RuihongQiu/DuoRec This method introduces unsupervised Dropout and supervised semantic augmentations in self-supervised learning for sequential recommendation.
  • FMLP-Rec: [^9] https://github.com/Woeee/FMLP-Rec Another most recent SR method adopting the full mlp layers architecture and filter methods on the self-attention module.